OpenStudy (clamin):

ONE MEDAL!! A card is randomly drawn from a regular deck of 52 cards.Determine the probability that the card drawn is a 2,4, or queen.

OpenStudy (anonymous):

probability of 2=1/13 probability of 4=1/13 probability of queen=1/13 total probability =1/13 *3=3/13

OpenStudy (clamin):

thats my answer too @apurba but my teacher said that its 12/52

OpenStudy (clamin):

im asking this question because i dont get it why its 12/52

OpenStudy (anonymous):

if u simplify 12/52 the ans is 3/13 both r the same...ur ans is correct

rvc (rvc):

probability 2=4/52 3=4/52 queen=4/52 total 12/52